# The Generic Temperature Response of Large Biochemical Networks

### v2

@@ II.8 Deviations from the generic quadratic response
Both our theory and computer simulations demonstrate that large networks show a generic Arrhenius plot with a quadratic form. We finally discuss deviations from this generic response. We start with small networks and demonstrate that they easily can be designed to give non-linear Arrhenius plots. We then discuss the case of linear chains, which shows that even large networks can deviate, because they contain only few spanning trees.
II.8.1 Small networks
@@
As an example for a small network, consider the $N_{v}=5$ network depicted in Fig. 13a. This network has two spanning trees rooted at vertex $4$ (see Fig. 13b:
@@
and four two-tree spanning forests, where one tree is rooted at vertex $4$ and the other contains vertex 0 (see Fig. 13c:
@@
II.8.2 Linear chain
The generic quadratic behaviour predicted here can be violated not only for small networks, but also for large networks which contain only few spanning trees. The simplest possible example is a linear chain of irreversible reactions, i.e., $k_{i,j}=k_{i}\delta_{i+1,j}$, as illustrated in Fig. (15)
Figure 15: A linear chain of $N$ reactions with reaction rates $k_{i,j}=k_{i}\delta_{i+1,j}$.
@@ II.8.2 Linear chain
Figure 16: Spanning forests of the linear chain. For $i\leq j$, the only corresponding spanning forest is the one without the $k_{j}$-edge. For $i>j$, there is no such tree.
@@ III Discussion
Being the result of evolution, the assumption that the activation energies can be described as independent from each other needs to be treated with care. But since there are many different reactions involved in a process such as the development of an embryo, some of which are purely biochemical (like the phosphorylations occuring in cell cycle control), some of which are more biophysical in nature (like the growth of microtubules), it seems reasonable to assume that most activation energies are indeed independent of each, especially when contained in different modules of the overall network. Furthermore, generalizations of the central limit theorem exist that allow for a weak degree of dependence, e.g. correlations decaying with distance in the network (see for example [87]). In the future, it might be interesting to experimentally determine all activation energies in a network of interest and then to predict the resulting temperature dependence using the formalism presented here.
@@
The theory developed so far is of course no longer valid above the melting temperature of the proteins. Denaturation slows down the reaction rates quite substantially and this is not described by the Arrhenius equation [85, 88]. This could explain the curvature of the Arrhenius plots seen at small values of the inverse temperature, in particular for temperatures of about $30^{\circ}\text{C}$ and above, where we observe that the high temperature data points (low inverse temperature) in Fig. 9-11 all lie below the quadratic curve. In the measurement by Powsner on the pupal development rates (Fig. 11B), the decrease is most notable, and it is not unlikely that this indicates that the denaturation threshold for a protein that is important for the developmental progression is surpassed at these temperatures. On the other hand, however, this observation also suggests that the larger part of the curve should not be affected by protein denaturation, and that our theory indeed captures the generic behaviour of this system. Future work could extend our theory to include the effects of protein denaturation, for example, by introducing it as an additional state in the network.
@@
The observation that the development rates can be well described by a quadratic fit in the Arrhenius plot does not seem to be limited to fly and frog as discussed above. Rombouts et al. [48] use it to fit the timing of the early cell cycles of Caenorhabditis elegans, Caenorhabditis briggsae, Danio rerio, Xenopus tropicalis and Xenopus laevis, and Aquilanti et al. [89] found that a quadratic dependence in the Arrhenius plot is a good description for data on the respiration rate of Camellia Japonica leaves.
@@ Appendix A Obtaining the MFPT by counting
Consider the one-step master equation with forward rates $k_{i}$ and backward rates $r_{i}$ as depicted in Fig. (17).
Figure 17: A one-step master equation from state $1$ to state $N+1$, the latter one being absorbing.
@@
The $\mathcal{F}_{[j,N+1]}^{1\rightarrow j}$ are the chains which start off towards the right and change their direction at $j$ and possibly also at a later vertex $m$. This is shown in Fig. 18.
Figure 18: Counting the spanning forests rooted at $N+1$ $\mathcal{F}_{[j,N+1]}^{1\rightarrow j}$ for the on-step master equation shown in Fig. (17).
Therefore, one gets:
@@
Figure 19: A simple KPR network on $N+1$ states consisting of a chain of forward rates $k_{i}$ and reset rates $\gamma_{i}$.
@@
The $\mathcal{F}_{[j,N+1]}^{1\rightarrow j}$ are the graphs with forward rates up to some vertex $j$ and either $k_{i}$ or $\gamma_{i}$ for any $i>j$. This is illustrated in Fig. 20.
Figure 20: Counting the $\mathcal{F}_{[j,N+1]}^{1\rightarrow j}$ in Fig. (19).
Therefore, one gets:
@@ Appendix B Fitting parameters for the figures in the main text
Appendix C Algorithm to compute the mean first-passage time in graph-theoretical framework
@@ 27 $F\leftarrow\text{compose}(T_{1,\text{oriented}},T_{2,\text{oriented}})$;
To illustrate that the sampling algorithm for the two-tree spanning forests works, it was applied to the one-step master equation from appendix A. The two-tree spanning forests were already characterized before. They consist of all choices of $j<N+1$ and $j\leq m<N-1$ corresponding to the forest with back rates from $m$ to $j$ and forward rates otherwise (compare also Fig. 18). So for a fixed number of back rates l=mjl=m-j, there are exactly NlN-l choices for $j$ and hence, NlN-l corresponding forests. In total, this means there are $\sum_{l=0}^{N}(N-l)=\frac{N(N+1)}{2}$ forests. So, if one draws uniformly a two-tree spanning forest, the probability $p(l)$ that it has $l$ back rates is given by:
@@
The convergence of the MFPT obtained via the algorithm to the correct value is shown in Fig. 26 (a). The exact value is given by Eq. (70).
@@
An advantage of this algorithm over simulating the stochastic dynamics of the network is that it allows a sample of spanning trees and two-tree spanning forests to be generated once, enabling the computation of the mean first passage time for different rates without regenerating the sample. This facilitates studying the impact of changing rates without the need to redo the entire simulation each time. To illustrate this, one sample of spanning forests of 50,176 spanning forests (from $1.000$ draws) was generated and the values of the MFPT were calculated for $k_{i}=1$ from $r_{i}=0$ to $r_{i}=0.9$ for all $i$ . The result is again compared to the exact solution given in Eq. (70). This is shown in Fig. 26 (b).
Appendix D Generating the total activation energy distributions
  • There must be a path from any vertex to the final vertex $N+1$.
  • The distance from vertex 0 to $N+1$ should be relatively large.
  1. 1.
  2. 2.
    • The $N$ vertices were randomly partitioned into $n$ non-empty sets.
  3. 3.
  4. 4.
    • For all cases, $n=N/10$ and $p=q=0.5$ were chosen.
The results of these simulations are presented in the main text.

### v3

@@ II.8 Deviations from the generic quadratic response
Both our theory and computer simulations demonstrate that large networks show a generic Arrhenius plot with a quadratic form. We now discuss deviations from this generic response that might arise because of our mathematical assumptions. We start with networks that have correlations between their activation energies and show that the generic response persists as long as such correlations are sufficiently dilute in large networks. We then discuss small networks and demonstrate that they easily can be designed to give non-linear Arrhenius plots. We finally discuss the case of linear chains, which shows that even large networks can deviate, because they contain only few spanning trees.
II.8.1 Correlations
Our theory assumes that activation energies on different edges are independent and that their summations over spanning trees and forests satisfy the conditions of the central limit theorem. In real biochemical networks, the master equation description contains repeated reaction channels due to multiplicity, so the corresponding activation energies are in fact correlated. As an instructive example, consider an autorepressive production of species $X$:
$\dot{X}=\frac{k}{1+(\frac{X}{C})^{n}},$ (39)
where $k$ is the production rate in the absence of $X$, $C$ is a reference concentration, and $n$ is the Hill coefficient. In a chemical master equation, the stochastic variable is the copy number $m$ of $X$ (for multiple species, it is the product space of all molecule counts). This yields a linear chain of reaction steps with rates $k_{m}=\frac{k}{1+(\frac{m}{c})^{n}}$, with $c$ being an appropriately rescaled constant (see Fig. 13).
Figure 13: Describing the autorepressive reaction in Eq. (39) as a master equation results in a linear chain of reactions where the rates all scale with the same reaction rate $k$.
Since the Arrhenius equation gives $k\sim e^{-\Delta\beta E_{A}}$, all these edges inherit the same activation energy $E_{A}$, because they are different instances of the same microscopic reaction. A more precise bookkeeping of total activation energy should therefore treat the energy of a spanning tree or forest as a weighted sum of distinct reaction types:
$E_{\mathcal{T}}=\sum_{k_{ij}}n_{ij}E_{ij},$ (40)
where $n_{ij}\in\mathbb{N}$ counts how many times that same reaction appears in the tree or forest. Although the state space of such master equations can be infinite in principle, in real biological systems the number of copies is effectively bounded due to resource limitations and hence it can be safely assumed that the first two moments of $n_{ij}$ stay finite. Note that the multiplicities $n_{ij}$ can themselves be treated as independent random variables and consequently $n_{ij}E_{ij}$ are independent random variables again. As long as there are still sufficiently many distinct, effectively independent reactions, Eq. (40) is still the sum of many independent random variables and the central limit theorem retains its validity.
This argument also applies to the possibility that the same network motif appears several times in the network, as known for example from the feedforward chain appearing in the developmental network of Caenorhabditis elegans [86]. Again this case could be remedied by introducing multiplicities $n_{ij}$. Only if the network was highly correlated throughout, would our generic prediction break down.
II.8.2 Small networks
@@
As an example for a small network, consider the $N_{v}=5$ network depicted in Fig. 14a. This network has two spanning trees rooted at vertex $4$ (see Fig. 14b):
@@
and four two-tree spanning forests, where one tree is rooted at vertex $4$ and the other contains vertex 0 (see Fig. 14c):
@@
II.8.3 Linear chain
The generic quadratic behaviour predicted here can be violated not only for small networks, but also for large networks which contain only few spanning trees. The simplest possible example is a linear chain of irreversible reactions, i.e., $k_{i,j}=k_{i}\delta_{i+1,j}$, as illustrated in Fig. 16.
@@ II.8.3 Linear chain
Figure 16: (a) A linear chain of $N$ reactions with reaction rates $k_{i,j}=k_{i}\delta_{i+1,j}$ (b) Spanning forests of the linear chain. For $i\leq j$, the only corresponding spanning forest is the one without the $k_{j}$-edge. For $i>j$, there is no such tree.
@@ III Discussion
Being the result of evolution, the assumption that the activation energies can be described as independent from each other needs to be treated with care. We first note that correlations arising from copies numbers in chemical master equations are not a fundamental limitation of our theory, because this effect can be described by multiplicities in the relevant sums. We also note that duplication of the same motif would not matter for the same reason. However, it could well be that two network motifs have been duplicated and then drifted apart during evolution. To date, no experimental evidence seems to be available to determine how strong such correlations might be in real developmental networks. However, since there are many different reactions involved in a process such as the development of an embryo, some of which are purely biochemical (like the phosphorylations occuring in cell cycle control), some of which are more biophysical in nature (like the growth of microtubules), it seems reasonable to assume that most activation energies are indeed independent of each other, especially when contained in different modules of the overall network. Furthermore, generalizations of the central limit theorem exist that allow for a weak degree of dependence, e.g. correlations decaying with distance in the network (see for example [88]). In the future, it might be interesting to experimentally determine all activation energies in a network of interest and then to predict the resulting temperature dependence using the formalism presented here.
@@
The theory developed so far is of course no longer valid above the melting temperature of the proteins. Denaturation substantially slows the reaction rates, and this is not described by the Arrhenius equation [85, 89]. This could explain the curvature of the Arrhenius plots seen at small values of the inverse temperature, in particular for temperatures of about $30^{\circ}\text{C}$ and above, where we observe that the high temperature data points (low inverse temperature) in Fig. 9-11 all lie below the quadratic curve. In the measurement by Powsner on the pupal development rates (Fig. 11b), the decrease is most notable, and it is not unlikely that this indicates that the denaturation threshold for a protein that is important for the developmental progression is surpassed at these temperatures. On the other hand, however, this observation also suggests that the larger part of the curve should not be affected by protein denaturation, and that our theory indeed captures the generic behavior of this system. Future work could extend our theory to include the effects of protein denaturation, for example, by introducing it as an additional state in the network.
@@
The observation that the development rates can be well described by a quadratic fit in the Arrhenius plot does not seem to be limited to fly and frog as discussed above. Rombouts et al. used it to fit the timing of the early cell cycles of Caenorhabditis elegans, Caenorhabditis briggsae, Danio rerio, Xenopus tropicalis and Xenopus laevis [48], and Aquilanti et al. found that a quadratic dependence in the Arrhenius plot is a good description for data on the respiration rate of Camellia Japonica leaves [90].
@@ Appendix A Obtaining the MFPT by counting
Consider the one-step master equation with forward rates $k_{i}$ and backward rates $r_{i}$ as depicted in Fig. 17.
Figure 17: (a) A one-step master equation from state $1$ to state $N+1$, the latter one being absorbing.
(b) Counting the two-tree spanning forests rooted at $N+1$ $\mathcal{F}_{[j,N+1]}^{1\rightarrow j}$ for this network.
@@
The $\mathcal{F}_{[j,N+1]}^{1\rightarrow j}$ are the chains which start off towards the right and change their direction at $j$ and possibly also at a later vertex $m$. This is shown in Fig. 17b. Therefore, one gets:
@@
Figure 18: (a) A simple KPR network on $N+1$ states consisting of a chain of forward rates $k_{i}$ and reset rates $\gamma_{i}$. (b) Counting the $\mathcal{F}_{[j,N+1]}^{1\rightarrow j}$ for this network.
@@
The $\mathcal{F}_{[j,N+1]}^{1\rightarrow j}$ are the graphs with forward rates up to some vertex $j$ and either $k_{i}$ or $\gamma_{i}$ for any $i>j$. This is illustrated in Fig. 18b. Therefore, one gets:
@@ Appendix B Fitting parameters for the figures in the main text
Appendix C Generating the total activation energy distributions
  • There must be a path from any vertex to the final vertex $N+1$.
  • The distance from vertex 0 to $N+1$ should be relatively large.
  1. 1.
  2. 2.
    • The $N$ vertices were randomly partitioned into $n$ non-empty sets.
  3. 3.
  4. 4.
    • For all cases, $n=N/10$ and $p=q=0.5$ were chosen.
The results of these simulations are presented in the main text.
Appendix D Algorithm to compute the mean first-passage time in graph-theoretical framework
@@ 27 $F\leftarrow\text{compose}(T_{1,\text{oriented}},T_{2,\text{oriented}})$;
To illustrate that the sampling algorithm for the two-tree spanning forests works, it was applied to the one-step master equation from appendix A. The two-tree spanning forests were already characterized before. They consist of all choices of $j<N+1$ and $j\leq m<N-1$ corresponding to the forest with back rates from $m$ to $j$ and forward rates otherwise (compare also Fig. 17b). So for a fixed number of back rates l=mjl=m-j, there are exactly NlN-l choices for $j$ and hence, NlN-l corresponding forests. In total, this means there are $\sum_{l=0}^{N}(N-l)=\frac{N(N+1)}{2}$ forests. So, if one draws uniformly a two-tree spanning forest, the probability $p(l)$ that it has $l$ back rates is given by:
@@
The convergence of the MFPT obtained via the algorithm to the correct value is shown in Fig. 24a. The exact value is given by Eq. (72).
@@
An advantage of this algorithm over simulating the stochastic dynamics of the network is that it allows a sample of spanning trees and two-tree spanning forests to be generated once, enabling the computation of the mean first-passage time for different rates without regenerating the sample. This facilitates studying the impact of changing rates without the need to redo the entire simulation each time. To illustrate this, one sample of spanning forests of 50,176 spanning forests (from $1.000$ draws) was generated and the values of the MFPT were calculated for $k_{i}=1$ from $r_{i}=0$ to $r_{i}=0.9$ for all $i$ . The result is again compared to the exact solution given in Eq. (72). This is shown in Fig. 24b.